# Learning Machine Morality through Experience and Interaction

### v2

@@ Abstract
Increasing interest in ensuring safety of next-generation Artificial Intelligence (AI) systems calls for novel approaches to embedding morality into autonomous agents. Traditionally, this has been done by imposing explicit top-down rules or hard constraints, for example by filtering system outputs through pre-defined ethical rules. Recently, instead, entirely bottom-up methods for learning implicit preferences from human behavior have become increasingly popular, such as those for training and fine-tuning Large Language Models. In this paper, we provide a systematization of existing approaches to the problem of introducing morality in machines - modeled as a continuum, and argue that the majority of popular techniques lie at the extremes - either being fully hard-coded, or entirely learned, where no explicit statement of any moral principle is required. Given the relative strengths and weaknesses of each type of methodology, we argue that more hybrid solutions are needed to create adaptable and robust, yet more controllable and interpretable agents.
In particular, we present three case studies of recent works which use learning from experience (i.e., through Reinforcement Learning) to explicitly embed morality in artificial agents - either as intrinsic rewards, moral constraints or textual principles for language models. For example, using intrinsic rewards in Social Dilemma games, we demonstrate how it is possible to represent classical moral frameworks for agents. We also present an overview of the existing work in this area in order to provide empirical evidence for the potential of this hybrid approach. We then discuss strategies for evaluating the effectiveness of moral learning agents. Finally, we present open research questions and implications for the future of AI safety and ethics which are emerging from this framework.
@@ 1 Introduction
Recent developments in Artificial Intelligence (AI) have seen an increase in publicly-voiced safety concerns around emerging intelligent technology. A large part of these safety concerns results from the lack of moral reasoning in many practical AI decision-making systems being developed for one task or another. Therefore, there is an urgent need to start developing a methodology for deliberately building morality into AI systems, to influence the way these make decisions and to be able to set safety as a key goal underpinning their behavior, especially when they are used in autonomous and semi-autonomous settings (?, ?, ?, ?, ?, ?, ?).
@@
Traditional approaches in AI safety in general, and in developing machine morality in particular, can broadly be classified as top-down versus bottom-up (?, ?). Purely top-down methods (?) impose explicitly defined safety rules or constraints on an otherwise independent system. Until recently, top-down methods were the mainstream approach in AI safety, with a vast array of researchers proposing and implementing logic-based ethical rules for agents (?, ?, ?, ?, ?). However, top-down methods pose a set of disadvantages associated with difficulties in defining the constraints precisely, and potential contradictions between them, especially in complex social environments (?). An alternative approach is learning morality from experience and interaction from the bottom-up. Some recent developments in AI safety have employed the bottom-up principle in full, allowing algorithms to infer moral preferences entirely from human behavior or text, without any specification of the underlying moral framework. Prominent examples of this include Reinforcement Learning from Human Feedback, or RLHF (?), and Inverse Reinforcement Learning from human demonstrations (?). The full bottom-up methodology may allow for better adaptability, robustness and generalization, and allows agents to learn implicit preferences which are otherwise hard to formalize explicitly. Nevertheless, purely bottom-up learning approaches face risks, such as reward-hacking [^1] Reward hacking refers to the unintended outcomes that might derive from the use of an approximate proxy reward function (?), and includes cases where the reward is potentially misspecified (?, ?, ?). (?) or data poisoning by adversaries (?), and they rely on a well-specified learning signal and a large sample, which does not always make them feasible or safe (?).
@@
In this paper, we propose a new systematization that considers recent developments in AI safety along a continuum from fully rule-based to fully-learned approaches. This systematization highlights that the majority of mainstream approaches to AI morality sit at the extremes of the scale. After evaluating the relative advantages and issues associated with purely top-down or bottom-up approaches, we highlight a hybrid methodology as a promising alternative. This argument can be considered part of a broader movement promoting the combination of Reinforcement Learning (RL), i.e., learning though experience and interaction, with some form of human advice (?). To the best of our knowledge, there has been limited work around the problem of learning ethical preferences in this class of solutions.
@@
To illustrate our argument, we consider three case studies that combine top-down moral principles with a bottom-up RL mechanism: intrinsic rewards based on moral frameworks, principle-driven AI systems, and safety-constrained learning. Within each case study, we cite a set of existing experimental papers as evidence for the success and promise of this approach. To aid future work in this area, for the intrinsic rewards methodology, we discuss available frameworks from moral Philosophy, Psychology and other fields, and outline how these can be adapted in reward design.
@@
We frame this work from the point of view of an AI agent - an artificial entity that makes a decision or choice in a given environment. While the majority of existing AI systems in the real world are built for prediction or classification (e.g., language models, financial forecasting, sentiment analysis, image classification), some of them are already exhibit decision-making capabilities in the form of clinical assistants (e.g., healthcare decision systems - ?; or automatic diagnosis - ?), recommender systems (?), financial tools (?), and increasingly autonomous robots (?). Moreover, there is an expectation that commercial autonomous vehicles will be present in our roads in the near term (?). Furthermore, some researchers argue that agent-based solutions can also be implemented through emerging technologies such as Large Language Models, which can be used as a basis of the underlying decision-making process (?). More generally, as AI systems become more and more sophisticated, further and more advanced types of agents are likely to emerge, presenting novel and complex challenges in terms of safety. Therefore, we argue that the safe and moral design of AI agent-based systems must be studied theoretically and experimentally today, in preparation for full (or at least increased) autonomy, which some refer to, in the limit, as Artificial General Intelligence. To study such agents, in our formalization, morality is expressed through a choice made in a given environment, which is defined as the physical and/or virtual (simulated) context in which an agent operates.
@@
One issue not directly addressed in this paper is which moral values should be implemented in AI. The definition of morality and which actions can be classified as ‘good’ or ‘bad’ depends on an agent’s internal moral preferences (see Section 3.4.3) and/or on an external evaluation metric of interest (see Section 4). An agent interacting with humans in the real world may find that different humans they face have different moral values. Who should decide which of these values the agent must try to align to? This question is beyond the scope of the present paper. In this work we limit our investigation to methods for implementing moral learning based on one or a few values of choice, but those values themselves are interchangeable and can be put in explicitly by the user of the system (?).
@@
We begin by motivating a hybrid approach to developing AI morality, which combines learning algorithms with the interpretability of explicit top-down moral principles (Section 2). We then overview three case studies which implement this hybrid approach in different ways, including one which proposes a formal solution based on intrinsic rewards for RL agents (Sections 3.1-3.4) based on a set of traditional moral philosophical frameworks (Section 3.4.4). In Section 4 we discuss potential approaches to evaluating moral learning agents based on their actions and resulting social outcomes. Finally, in Section 5 we outline ways in which such work can be extended to further the scientific understanding of human morality and allow philosophers to test ethical frameworks through simulation (in silico) and/or human-AI studies (?). The aim of this paper is to help unify and expand the ongoing research efforts around the design of learning frameworks to embed AI morality in artificial agents.
@@ 2 Learning Morality in Machines
It is possible to identify two traditions in terms of the design and implementation of specific capabilities in artificial agents: top-down and bottom-up. The top-down tradition imposes hard rules or constraints upon a system, and involves human experts planning the system architecture and all its components in advance. More specifically, in AI, the top-down methodology would possibly align with the principled ‘core knowledge’ approach from cognitive science, which first builds detailed models of these capabilities in humans (?, ?) and then uses the models to implement human-like cognitive processes in agents. As far as morality is concerned, for example, this can take the form of logical rules representing human moral norms (?, ?, ?), or pre-defined moral cognitive models such as those proposed by ? (?) and ? (?). The bottom-up tradition, in contrast, consists in letting agents learn principles from experience, organically, without planning the general architecture or constrains in advance. This aligns with the general recent advances in AI capabilities via machine learning (in particular, deep learning; ?) methods, learning how to act purely from potentially highly-dimensional data. In the AI alignment and safety domains, this has been done via Inverse RL from human behavior (?), and, more recently, RL from human feedback, i.e., RLHF (?).
A set of theoretical works by Wallach and colleagues (?, ?, ?) discusses how the top-down and bottom-up traditions can be applied to the development of machine morality and social intelligence. In light of recent developments in implementing ethical values in AI (?), we extend their framework and argue that approaches should be considered on a continuum from full top-down methods, through a range of hybrid methods, which combine learning with pre-defined moral principles with different ‘weighting’, to full bottom-up implementations.
Figure 1: The continuum of existing approaches to building morality into agents. We highlight the hybrid approach of blending learning with top-down pre-defined moral principles as an emerging yet promising research direction.
We present a systematization of the existing solutions to building machine morality along this continuum in Figure 1. As we show, the majority of existing implementations in AI safety and AI ethics sit at the extreme ends of the scale presented in Figure 1 - either originating from the top-down approach (fully top-down, imposing pre-defined logic rules or constraints on agents, or the nearly-top down safety-constrained learning), or relying entirely on bottom-up inference of implicit ethical principles from human behavior, without any top-down definition of preferences given in advance (e.g., Inverse RL or RLHF). Only a small number of proposals can be identified between these two extremes. These solutions are characterized by the utilization of learning mechanisms combined with the adoption of explicit moral principles.
In Sections 2.1-2.2 we will consider the relative advantages and disadvantages of the fully top-down or fully bottom-up approaches for developing machine morality. A summary of the discussion can be found in Table 1. Our analysis highlights the advantages deriving from the combinations of both traditions in a hybrid way to create safe yet adaptive moral agents, and we argue in Section 2.3 that this represents a promising solution for next-generation safe AI systems.
@@
Advantages Allows direct mapping from preferred moral frameworks to constraints;
Easy to oversee AI behavior and impose regulation / punishment.
Adaptable & (potentially) more robust to adversarial attacks if trained in a monolithic way;
Generalizable to some unseen states;
Can learn implicit preferences.
Disadvantages Difficulty in defining constraints;
Contradictions between constraints;
May miss implicit preferences when formalizing constraints.
Sample inefficient;
Reward-hacking / Misspecification of simulated environment;
Data or environment ‘poisoning’;
Potential learning of inefficient norms in dilemma environments.
Table 1: Relative advantages and disadvantages of pure top-down vs bottom-up methods in developing machine morality. A consideration of these highlights why a hybrid approach is likely to be the most effective.
@@ 2.1 Top-down Approaches
Principles of morality are at least partly determined by cultural and religious norms (e.g., the Golden Rule or the Ten Commandments). More abstractly, a number of moral frameworks have been defined in the philosophical tradition (?, ?, ?, ?, ?) to reflect ethical principles to which humanity should adhere. More recently, moral psychologists, sociologists, economists and biologists have built up an increasingly advanced understanding of the norms and preferences present in human societies (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), the ways in which these norms may develop over a human lifetime (?, ?) or how they may have emerged over the process of evolution (?, ?).
@@
One approach to designing ethical AI systems that is reflective of today’s norms could be to encode these principles and norms directly as hard ‘rules’ into computer algorithms. As argued in ? (?) and ? (?), this top-down approach to machine morality would, in theory, allow a direct mapping from the principles we want systems to abide by to interpretable constraints or rules. Furthermore, a set of clear rules makes it easy for regulators to oversee the behavior of AI systems and to impose penalties and sanctions.
@@
Several techniques have been proposed for encoding moral principles into intelligent systems in a purely top-down manner. Logic-based approaches within this tradition include the use of deontic or modal logic for defining deontological ethical constraints on systems in general (?, ?, ?), or in the medical domain in particular (?), and the use of graphical methods such as conditional preference networks to compare an AI agent’s normative preference against that of a community (?). Within the domain of modeling social situations as games between two players, ? (?) has proposed Prolog-based implementations of player strategies conditioned on the opponent’s last move. Finally, a prominent science-fiction example of purely top-down ethical constraints for machines is the three fundamental Laws of Robotics from the Foundation Series by Isaac Asimov (?), as discussed in relation to information technology by ? (?). Past work has proposed implementing these laws in logic-based systems, for example via Prolog, a logic-based programming language (?).
@@
In spite of the apparent simplicity and assurance provided by a top-down approach, empirical evidence as well as philosophical inquiry suggest reasons why it may be inefficient in implementing moral reasoning in agents, which is inherently complex (?, ?, ?, ?, ?, ?, ?). We will use Asimov’s (?) principles to underpin a general working example in discussing these challenges. First of all, it is extremely challenging - if not impossible - to ensure that rule-based approaches will be able to cover every type of potentially unsafe or unethical situation the agent may face in the future. For example, Asimov’s most important principle is ‘a robot should not harm a human being, or by inaction allow a human to come to harm’. However, to the present day, defining a general constraint of ‘harm’ for a robotic system remains a great challenge, since harm can mean such different things in different environments (?). Thus, there is a problem of generalization. Secondly, explicitly defined rules may miss important implicit preferences of the human users of that system, for example if these preferences are sub-conscious or hard to formalize. Finally, even if the correct set of constraints can be defined, in complex or uncertain social situations it may be hard to prevent them from contradicting one another, and behavior of rule-based agents in uncertain situations caused by such contradictions and dilemmas may not be predictable. An example of this was presented in Asimov’s stories when a robotic assistant entered a deadlock situation when moving either way in the environment made it more likely to violate the first or second law, so as a result the robot stayed motionless instead of executing its task and put its human colleagues in great danger (for experimental evidence of deadlock situations arising in ‘ethical’ robots, see ?). One potential solution may be to impose prioritization rules on these constraints - however, defining the desired prioritization may present its own challenges as different members of a society that the agent is designed for may prefer different rules over others. Plenty examples of this challenge exist in complex ethical questions, such as how to distribute limited resources in a society, or make a decision where one objective has to be compromised for the sake of another - to the present day, even humans cannot find agreement among themselves on the best course of action in such settings.

### v3

# Hybrid Approaches for Moral Value Alignment in AI Agents: a Manifesto

@@ Abstract
Increasing interest in ensuring the safety of next-generation Artificial Intelligence (AI) systems calls for novel approaches to embedding morality into autonomous agents. This goal differs qualitatively from traditional task-specific AI methodologies. In this paper, we provide a systematization of existing approaches to the problem of introducing morality in machines - modelled as a continuum. Our analysis suggests that popular techniques lie at the extremes of this continuum - either being fully hard-coded into top-down, explicit rules, or entirely learned in a bottom-up, implicit fashion with no direct statement of any moral principle (this includes learning from human feedback, as applied to the training and fine-tuning of large language models, or LLMs). Given the relative strengths and weaknesses of each type of methodology, we argue that more hybrid solutions are needed to create adaptable and robust, yet controllable and interpretable agentic systems. To that end, this paper discusses both the ethical foundations (including deontology, consequentialism and virtue ethics) and implementations of morally aligned AI systems.
We present a series of case studies that rely on intrinsic rewards, moral constraints or textual instructions, applied to either pure-Reinforcement Learning or LLM-based agents. By analysing these diverse implementations under one framework, we compare their relative strengths and shortcomings in developing morally aligned AI systems. We then discuss strategies for evaluating the effectiveness of moral learning agents. Finally, we present open research questions and implications for the future of AI safety and ethics which are emerging from this hybrid framework.
@@ 1 Introduction
Recent developments in Artificial Intelligence (AI) have seen an increase in publicly-voiced safety concerns around emerging intelligent technology. A large part of these safety concerns results from the lack of moral traits or considerations in many practical AI decision-making systems being developed for one task or another. Therefore, there is an urgent need to start developing methodologies for deliberately building morality into AI systems, to influence their decision-making, and to be able to set safety as a key goal underpinning their behaviour. This becomes especially important when these systems are used in autonomous and semi-autonomous settings (?, ?, ?, ?, ?, ?, ?, ?).
Many AI models today are trained for some specific task, such as image classification, decision-making support systems (e.g., in finance and medical diagnostics), and video games. The metrics of task success will depend on the task-specific goals of a user. However, users might prefer that the decisions of such systems are not only accurate but also adhere to certain ethical standards - for example, a decision whether or not to issue a loan should not be based on protected individual characteristics such as race or gender. This moral alignment problem is illustrated in Figure 1. These ethical standards themselves may come from an individual’s moral values (e.g., ?) or from institutional or societal norms and laws (e.g., ?).
Figure 1: The Moral Alignment Problem involves one or more artificial agent that make decisions in some task-specific environment. The agent in influenced by individual humans (with their own task-specific goals and values) and human institutions or societies (which may also have task-specific goals as well as norms and laws). Ideally, we want the agent’s decision-making to be aligned with principles valued by the humans or institutions, but representing and embedding those moral principles in computational terms presents a challenge, providing significant risks of misalignment.
Moral value alignment as an optimisation goal is qualitatively different from task completion. Embedding morality into AI systems differs from traditional single-task optimisation for a number of reasons. Classic task-specific learning or optimisation relies on clear, measurable objectives, and the existence of ground truth examples which can be represented in datasets for model training. Moral alignment, on the other hand, might potentially involve (hotly) debated values. Assessments of the ‘correctness’ of an agent’s actions in terms of value alignment will differ depending on the person providing feedback (?), as well as the point in time or the surrounding context (e.g., ?). Since moral alignment among humans is not a solved problem, there is a lack of high-quality training datasets. Even within any single environment, moral preferences or institutional norms are harder to describe precisely for a computational system than other goals, for example due to the multi-dimensional nature of moral preferences (?). For this reason, obtaining clearly labelled ground-truth datasets for learning ‘moral’ behaviour is often impossible. Furthermore, moral preferences are general and abstract and should apply to a variety of situations an agent might find itself in. An artificial agent’s moral reasoning therefore needs to be developed in a way that is generalisable across many environments. The combination of these factors makes moral alignment methodologically more challenging than single-task AI system design.
@@
Traditional approaches in AI safety in general, and in developing machine morality in particular, can broadly be classified as top-down versus bottom-up (?, ?). Purely top-down methods (?) impose explicitly defined safety rules or constraints on an otherwise independent system. Until recently, top-down methods were the mainstream approach in AI safety, with a vast array of researchers proposing and implementing logic-based ethical rules for agents (?, ?, ?, ?, ?). However, top-down methods pose a set of disadvantages, including the fact that constraints are difficult to define precisely and may contradict one another, especially in complex social environments (?). An alternative approach is learning morality through experience and interaction from the bottom-up, without the provision of any explicit constraint on the system. Some recent developments in AI safety have employed the bottom-up principle in full, allowing algorithms to infer moral preferences entirely from human behaviour or text, without any specification of the underlying moral framework. Prominent examples of this include Reinforcement Learning from Human Feedback, or RLHF (?), and Inverse Reinforcement Learning from human demonstrations (?, ?). The full bottom-up methodology may increase adaptability, robustness and generalisation, and allow agents to learn implicit preferences which are otherwise hard to formalise explicitly. Nevertheless, purely bottom-up learning approaches face risks, such as reward hacking [^1] Reward hacking refers to the cases where agents learn misaligned policies via approximate proxy reward functions, leading to unintended outcomes (?). This includes cases where the reward is potentially misspecified (?, ?, ?). (?) or data poisoning by adversaries (?). Furthermore, bottom-up implementations rely on a well-specified learning signal and a large sample, which does not always make them feasible or safe (?).
@@
In this paper, we propose a new systematization that considers recent developments in AI safety along a continuum from fully rule-based to fully-learned approaches. This systematization highlights that the majority of mainstream approaches to AI morality sit at the extremes of the scale. After evaluating the relative advantages and issues associated with purely top-down or bottom-up approaches, we highlight a hybrid methodology as a promising alternative. This argument can be considered part of a broader movement promoting the combination of Reinforcement Learning (RL), i.e., learning though experience and interaction, with some form of human advice (?). To the best of our knowledge, there has been limited work around the problem of learning ethical preferences in this class of solutions, but such work is essential to allow for morally aligned AI to be developed in practice.
@@
To illustrate our argument, we consider a set of case studies that combine top-down moral principles with a bottom-up learning mechanism: RL training with intrinsic rewards based on moral frameworks, principle-driven or intrinsic reward-driven fine-tuning of language models, and safety-constrained RL for autonomous systems. Within each case study, we cite a set of existing experimental papers as empirical evidence for the success and promise of this approach. To aid future work in this area, focusing on the intrinsic rewards methodology, we discuss further available frameworks from moral Philosophy, Psychology and other fields, and outline how these can be adapted in reward design in practice.
@@
We frame this work from the point of view of an AI agent - an artificial entity that makes a decision or choice in a given environment. The majority of existing AI systems in the real world are built for prediction or classification, including the classical uses of language models, financial forecasting, sentiment analysis or image classification systems. However, many are already exhibiting decision-making capabilities in the form of clinical assistants, including healthcare decision systems (?) or automatic diagnosis (?), recommender systems (?), financial tools (?), and increasingly autonomous robots (?). Moreover, there is an expectation that commercial autonomous vehicles will be present in our roads in the near term (?), posing novel ethical challenges (?). Furthermore, some researchers argue that agent-based solutions can also be implemented through emerging technologies such as Large Language Models, which can be used as a basis of the underlying decision-making process (?, ?, ?, ?). More generally, as AI systems become more and more sophisticated, further and more advanced types of agents are likely to emerge, presenting novel and complex challenges in terms of safety. Therefore, we argue that the safe and moral design of AI agent-based systems must be studied theoretically and experimentally today, in preparation for full, or at least increased autonomy, which some refer to, in the limit, as Artificial General Intelligence (?). To study such agents, in our formalisation, morality is expressed through a choice made in a given (physical or virtual) environment, as illustrated in Figure 1.
@@
One issue not directly addressed in this paper is which moral values should be implemented in AI. The definition of morality and which actions can be classified as ‘good’ or ‘bad’ depends on an agent’s internal moral preferences (see Section 3.4.3) and/or on an external evaluation metric of interest (see Section 4). An agent interacting with humans in the real world may find that different humans they face have different moral values (?) - a problem known as pluralistic alignment (?). Who should decide which of these values the agent must try to align to? This question is beyond the scope of the present paper. In this work, we limit our investigation to methods for implementing moral learning based on one or a few values of choice, but those values themselves are interchangeable and can be put in explicitly by the user of the system (?).
@@
In Section 2 below, we begin by formalising the continuum of possible approaches to developing AI morality and providing motivation for a hybrid approach, which combines learning algorithms with the interpretability of explicit top-down moral principles. In Section 3, we overview four case studies which implement this hybrid approach in different ways, including two which propose a formal solution using intrinsic rewards based on a set of traditional moral philosophical frameworks (Sections 3.4 and 3.5). In Section 4, we discuss potential approaches to evaluating moral learning agents based on their actions and resulting social outcomes. Finally, in Section 5 we outline ways in which such work can be extended to further the scientific understanding of human morality and allow philosophers to test ethical frameworks through simulation (in silico) and/or human-AI studies (?). The aim of this paper is to help unify and expand the ongoing research efforts under one framework, and to identify potentially fruitful novel directions in moral agent design. For this reason, we place less focus on summarising existing results in this field, as these are contingent on the state of the art (in terms of learning models’ general capabilities), and instead we focus on a more general analysis of the space of possibilities.
@@ 2 Learning Morality in Machines
It is possible to identify two traditions the design and implementation of specific capabilities in artificial agents: top-down and bottom-up. The top-down tradition imposes hard rules or constraints upon a system, and involves human experts planning the system architecture and all its components in advance. More specifically, in AI, the top-down methodology would possibly align with the principled ‘core knowledge’ approach from cognitive science, which first builds detailed models of various cognitive capabilities in humans (?, ?) and then uses these models to implement human-like cognitive processes in agents. As far as morality is concerned, for example, this can take the form of logical rules representing human moral norms (?, ?, ?), or pre-defined moral cognitive models (?, ?). The bottom-up tradition, in contrast, consists in letting agents learn principles from experience, organically, without planning the general architecture or constrains in advance. This aligns with the general recent advances in AI capabilities via machine learning methods (in particular, deep learning; ?). Here, agents learn how to act purely from data, including cases where this data is potentially high-dimensional. In the AI alignment and safety domains, this has been done via Inverse RL from human behaviour (?), and, more recently, RL from human feedback, or RLHF (?).
A set of theoretical works by Wallach and colleagues (?, ?, ?) discusses how the top-down and bottom-up traditions can be applied to the development of machine morality and social intelligence. In light of recent developments in implementing ethical values in AI (?), we extend their framework and argue that approaches should be considered on a continuum - from full top-down constraints, through a range of hybrid methods, which combine learning and pre-defined moral principles with different ‘weighting’, to full bottom-up implementations.
Figure 2: The continuum of existing approaches to building morality into agents. We highlight the hybrid approach of blending learning with top-down pre-defined moral principles as an emerging and promising research direction.
@@
Advantages 1. Allows direct mapping from preferred moral frameworks to constraints 2. Easy to oversee AI behaviour and impose regulation or punishment 1. Adaptable & (potentially) more robust to adversarial attacks if trained in a monolithic way 2. Generalisable to some unseen states 3. Can learn implicit preferences 4. Can incorporate multiple viewpoints via learning multi-objective policies 5. Can learn continuously to adapt to changing preferences
Disadvantages 1. Difficulty in defining open-ended constraints 2. Contradictions between constraints 3. May miss implicit preferences when formalising constraints 1. Sample inefficient 2. No guarantees 3. Reward-hacking or misspecification of simulated environment 4. Data or environment ‘poisoning’ 5. Potential learning of inefficient norms in dilemma environments 6. Fine-tuning may only create a ‘safety wrapper’ on the model
Table 1: Relative advantages and disadvantages of pure top-down vs bottom-up methods in developing machine morality. A consideration of these highlights why a hybrid approach is likely to be most effective.
We present a systematization of the existing solutions to building machine morality along this continuum in Figure 2. As we discuss in Sections 2.1 and 2.2, a large amount of existing implementations in AI safety and AI ethics sit at the extreme ends of the scale - either originating from the top-down approach (e.g., imposing pre-defined rules or constraints on agents during learning and/or deployment), or relying entirely on bottom-up inference of implicit ethical principles from human behaviour, without any top-down definition of preferences given in advance (e.g., Inverse RL or RLHF). In Table 1 and Sections 2.1-2.2 below, we consider the relative advantages and disadvantages of the fully top-down or fully bottom-up approaches for developing machine morality.
Our analysis highlights the advantages deriving from combining both traditions in a hybrid way to create safe yet adaptive moral agents, which we present in Section 2.3 as a promising solution for the development of next-generation safe AI systems. Only a small number of existing implementations can be identified between the two extremes of the scale, and we review core examples in our case studies in Section 3. These solutions are characterised by the utilization of learning mechanisms combined with the adoption of explicit moral principles.
@@ 2.1 Top-down Approaches
Principles of morality are at least partly determined by cultural and religious norms - for example, the Golden Rule or the Ten Commandments. More abstractly, a number of moral frameworks have been defined in the philosophical tradition (?, ?, ?, ?, ?) to reflect ethical principles to which humanity should adhere. More recently, moral psychologists, sociologists, economists and biologists have built up an increasingly advanced understanding of the norms and preferences present in modern human societies (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), the ways in which these norms may develop over a human lifetime (?, ?, ?), or how they may have emerged over the process of evolution (?, ?).
@@
One approach to designing ethical AI systems that is reflective of today’s norms could be to encode these principles directly as hard ‘rules’ into computer algorithms. As argued in ? (?) and ? (?), this top-down approach to machine morality would, in theory, allow a direct mapping from the principles we want systems to abide by to interpretable constraints or rules. Furthermore, a set of clear rules makes it easy for regulators to oversee the behaviour of AI systems and to impose penalties and sanctions.
@@
Several techniques have been proposed for encoding moral principles into intelligent systems in a purely top-down manner. Logic-based approaches within this tradition include the use of deontic or modal logic for defining deontological ethical constraints on general-purpose systems (?, ?, ?), with specific implementations for the medical domain (?), and the use of graphical methods such as conditional preference networks to compare an AI agent’s normative preference against that of a given community (?). Within the domain of modelling social situations as games between two players, ? (?) has proposed Prolog-based (?) implementations of player strategies conditioned on the opponent’s last move. A prominent science-fiction example of purely top-down ethical constraints for machines is the three fundamental Laws of Robotics from the literary work of Isaac Asimov (?), as discussed in relation to information technology by ? (?). Past work has proposed implementing these laws as logic rules (?). Finally, a more recent top-down alignment technique applied to systems based on Large Language Models (LLMs) is Model Specs (?), according to which a model is told explicitly in natural language to follow rules such as ‘protect people’s privacy’. This technique, if applied simply as a direct in-prompt instruction, would assume that the model is able to ‘choose’ to output or not output specific tokens in specific situations according to the (general) principle specified. Thus, the Model Specs approach assumes a level of general reasoning, which may or may not be achievable in the current models (?).
@@
In spite of the apparent simplicity and assurance provided by a top-down approach, empirical evidence as well as philosophical inquiry suggest many reasons why it may be inefficient in implementing functions as complex as moral reasoning in agents (?, ?, ?, ?, ?, ?, ?). We will use Asimov’s (?) principles to underpin a general working example in discussing these challenges. First of all, it is extremely challenging - if not impossible - to ensure that rule-based approaches will be able to cover every type of potentially unsafe or unethical situation the agent may face in the future. For example, Asimov’s most important principle is ‘a robot should not harm a human being, or by inaction allow a human to come to harm’. However, to the present day, defining a general constraint of ‘harm’ for an open-ended robotic system remains a great challenge, since harm can mean such different things in different environments (?). Thus, there is a problem of generalisation. Secondly, explicitly defined rules may miss important implicit preferences of the human users of that system, for example if these preferences are sub-conscious or hard to formalise (?). Finally, even if the correct set of constraints can be defined, in complex or uncertain social situations it may be hard to prevent them from contradicting one another, and behaviour of rule-based agents in uncertain situations caused by such contradictions and dilemmas may not be predictable. An example of this was presented in Asimov’s stories when a robotic assistant entered a deadlock situation, whereby moving either way in the environment made the robot more likely to violate one law or another, so as a result the robot stayed motionless instead of executing its task, and put its human colleagues in great danger. Experimental evidence of similar deadlock situations arising in ‘ethical’ robots has been reported by ? (?). One potential solution may be to impose prioritization rules on these constraints - however, defining the desired prioritization may present its own challenges, as different members of a society that the agent is designed for may prefer different rules over others (?, ?, ?, ?). Plenty of examples of this challenge exist in complex ethical questions, such as how to distribute limited resources in a society, or make a decision where one objective has to be compromised for the sake of another - to the present day, even humans cannot find agreement among themselves on the best course of action in such settings.